Im not sure what category this goes under. but i would like to know how to answer these types of problems including:
cuberoot 5/16
cuberoot 4/17
things like that
: Im not sure what category this goes under. but i would like to know how to answer these types of problems including:
cuberoot 5/16
cuberoot 4/17
things like that

Answer by stanbon(19020)

cuberoot (5/16)
Multiply numerator and denominator by 4 to get:
cubert(20/64) = [cubert(20)] / [4] = (1/4)cuberoot(20)
----------------------
cuberoot 4/17
Multiply numerator and denominator by 17^2 to get:
[cuberoot[4*17^2] / 17 = (1/17)cuberoot(1156)
=================
Cheers,
Stan H.